# Kelix Technique Report: Closing the Understanding Gap of Discrete Tokens in Unified Multimodal Models

### v2

@@ Abstract
Autoregressive large language models (LLMs) scale well by expressing diverse tasks as sequences of discrete natural-language tokens and training with next-token prediction, which unifies comprehension and generation under self-supervision. Extending this paradigm to multimodal data requires a shared, discrete representation across modalities. However, most vision-language models (VLMs) still rely on a hybrid interface: discrete text tokens paired with continuous Vision Transformer (ViT) features. Because supervision is largely text-driven, these models are often biased toward understanding and cannot fully leverage large-scale self-supervised learning on non-text data. Recent work has explored discrete visual tokenization to enable fully autoregressive multimodal modeling, showing promising progress toward unified understanding and generation. Yet existing discrete vision tokens frequently lose information due to limited code capacity, resulting in noticeably weaker understanding than continuous-feature VLMs. We present Kelix, a fully discrete autoregressive unified model that closes the understanding gap between discrete and continuous visual representations. At its core is a novel multi-token vision tokenization mechanism that increases the information capacity of discrete codes, enabling discrete-token models to achieve understanding performance comparable to their continuous-feature counterparts. Extensive experiments demonstrate state-of-the-art results on most multimodal understanding and image generation benchmarks among comparable-scale unified models. In particular, Kelix achieves significant improvements over prior discrete approaches on understanding tasks, especially on text-rich benchmarks, where discrete methods have been weakest, surpassing the previous best discrete model by over 23% on OCRBench and closely approaching continuous baselines. This demonstrates that discrete representations are no longer the bottleneck of multimodal understanding, offering new insights for the future development of discrete unified models.
@@ 1 Introduction
In this paper, we propose Kelix [^1] The name Kelix is a portmanteau of Kuaishou and helix: just as the DNA double helix encodes the full complexity of life using only four discrete nucleotide bases (A, T, C, G), Kelix encodes rich visual semantics using discrete tokens., a fully discrete, LLM-centric unified model that closes this understanding gap. At its core is a multi-token vision tokenization scheme that expands the coding space of discrete visual representations, substantially reducing information loss. The tokenizer is learned end-to-end with the LLM, allowing quantization to adapt to the downstream objective. We further adopt a next-block prediction objective that models grouped discrete token blocks, mitigating context-length explosion while preserving multimodal information. Extensive experiments demonstrate state-of-the-art results on major understanding and generation benchmarks among comparable-scale unified models.
@@ 2 Design Philosophy
Our approach. We address this bottleneck with a multi-token discretization scheme that substantially increases the information capacity of discrete vision tokens, yielding understanding on par with continuous-feature VLMs while remaining fully discrete and autoregressive. For the visual modality, we use a ViT-based encoder as the Tokenizer and a diffusion-based decoder as the Detokenizer, letting each component focus on its strength: the Tokenizer aligns visual signals into the shared semantic space, the LLM handles reasoning, and the Detokenizer renders high-quality images.
@@ 3 Architecture
This section begins by introducing the discrete image tokenizer, detailing how to quantize the ViT-produced patch embeddings via a multi-token quantization mechanism. We then introduce a next-block-prediction training paradigm for the unified LLM backbone, which compresses the increased vision tokens from our multi-token tokenizer into manageable blocks, thereby preserving training and inference efficiency comparable to single-token methods. Finally, we describe the diffusion-based de-tokenizer that renders high-quality images conditioned on the LLM-generated image tokens.
@@ 3.1 Discrete Image Tokenizer
Multi-Token Quantization. To address this bottleneck, we decompose each continuous patch feature into $N$ parallel subspaces via dedicated projection matrices, and quantize each subspace independently with its own sub-codebook. The combinatorial coding capacity thus grows exponentially with $N$, substantially narrowing the gap with continuous representations. To keep the context length unchanged for the LLM encoder, we apply sum pooling to aggregate the $N$ quantized sub-vectors back into a single composite token on the encoder side, while the decoder retains the full $N$-way discrete codes for autoregressive prediction. This asymmetric design balances representational expressivity with computational efficiency.
@@
Architecture and Training Objective. Kelix-Tok comprises two core components: (1) a NaViT-style vision encoder, initialized from Keye-VL and supporting native-resolution input, and (2) the multi-token quantization layer described above, which decomposes each continuous patch feature into $N$ parallel discrete codes. To align the resulting discrete visual representations with the LLM text space, we attach a pre-trained lightweight language model (600M parameters) as an auxiliary decoder during training. The overall objective jointly optimizes a next-token prediction loss and a quantization loss:
@@ 3.2 Unified Model Architecture
As discussed in Section 3.1, our multi-token quantization produces $N$ discrete codes per patch, substantially increasing the number of vision tokens compared to the original ViT output. To handle this without inflating the context length, we adopt a unified autoregressive Transformer architecture built around a Next-Block Prediction (NBP) paradigm.
@@ 5.2 Image Understanding Performance
By contrast, Kelix stands as the first discrete-tokenization unified model to achieve OCRBench performace on par with that of understanding-only VLMs. Specifically, Kelix attains an OCRBench score of 86.7, which is nearly identical to the 86.4 achieved by VLM Qwen2.5-VL (7B), while outperforming the second-best discrete-tokenization baseline X-Omni (7B) by a substantial 16.3 points, corresponding to a relative performance improvement of 23%. The result validates that Kelix’s multi-token quantization mechanism effectively bridges the representational gap between continuous visual embeddings and discrete tokens, demonstrating that discrete visual representations are no longer a fundamental performance bottleneck for multimodal understanding even for text-rich tasks that demand fine-grained semantic preservation.
@@ 6.1 Ablation Study on Kelix-Tok
Baseline Comparison. The results in Table 6(a) reveals a critical trade-off among different visual representation schemes. We observe that, compared to the Continuous Baseline which represents a performance upper bound, the Discrete Baseline introduces a significant performance gap. This is particularly pronounced on Text-rich tasks, where its performance drops to approximately 85% of the continuous baseline (66.5 vs. 78.0 Avg.) due to the information bottleneck. Our Kelix-Tok, leveraging a multi-token combinatorial representation, successfully achieves discretization while attaining performance comparable to its continuous counterpart.
@@
Number of Sub-Tokens. We investigate the impact of varying the number of sub-tokens $N$ by conducting experiments with $N=1,2,4,8,16$ alongside a continuous baseline, with results presented in Fig. 7. As shown in Subfigure. 7(a), the LM loss consistently decreases with increasing $N$, and the loss curves for $N\geq 8$ converge closely to the continuous baseline, indicating that sufficient discrete tokens can effectively compress continuous visual information without substantial loss. We further evaluate our training checkpoints on the OCRBench benchmark, which are highly sensitive to model performance. Subfigure. 7(b) demonstrates that OCRBench scores consistently increase with $N$. Upon training convergence, the performance of the continuous baseline is nearly matched by the top two discrete configurations ($N=8$ and $N=16$), which further validates the robustness of our multi-token based Kelix.
@@ 7.2 Unified Multimodal Models
Our Position. Our proposed Kelix follows the unified architecture route with high-level discrete semantics, but introduces a novel multi-token quantization mechanism that expands the representational capacity of discrete tokens without increasing the LLM’s sequence length. This directly addresses the information loss bottleneck of discrete tokenization, aiming to close the understanding gap between discrete and continuous representations and unlock the full potential of unified multimodal modeling.

### v3

# Kelix Technical Report: Closing the Understanding Gap of Discrete Tokens in Unified Multimodal Models

@@ Abstract
Autoregressive large language models (LLMs) scale well by expressing diverse tasks as sequences of discrete natural-language tokens and training with next-token prediction, which unifies comprehension and generation under self-supervision. Extending this paradigm to multimodal data requires a shared, discrete representation across modalities. However, most vision-language models (VLMs) still rely on a hybrid interface: discrete text tokens paired with continuous Vision Transformer (ViT) features. Because supervision is largely text-driven, these models are often biased toward understanding and cannot fully leverage large-scale self-supervised learning on non-text data. Recent work has explored discrete visual tokenization to enable fully autoregressive multimodal modeling, showing promising progress toward unified understanding and generation. Yet existing discrete vision tokens frequently lose information due to limited code capacity, resulting in noticeably weaker understanding than continuous-feature VLMs. We present Kelix, a fully discrete autoregressive unified model that closes the understanding gap between discrete and continuous visual representations. At its core is a product-quantization-based vision tokenization mechanism that increases the information capacity of discrete codes, enabling discrete-token models to achieve understanding performance comparable to their continuous-feature counterparts. Extensive experiments demonstrate state-of-the-art results on most multimodal understanding and image generation benchmarks among comparable-scale unified models. In particular, Kelix achieves significant improvements over prior discrete approaches on understanding tasks, especially on text-rich benchmarks, where discrete methods have been weakest, surpassing the previous best discrete model by over 23% on OCRBench and closely approaching continuous baselines. This demonstrates that discrete representations are no longer the bottleneck of multimodal understanding, offering new insights for the future development of discrete unified models.
@@ 1 Introduction
In this paper, we propose Kelix [^1] The name Kelix is a portmanteau of Kuaishou and helix: just as the DNA double helix encodes the full complexity of life using only four discrete nucleotide bases (A, T, C, G), Kelix encodes rich visual semantics using discrete tokens., a fully discrete, LLM-centric unified model that closes this understanding gap. At its core is a product-quantization-based vision tokenization scheme that expands the coding space of discrete visual representations, substantially reducing information loss. The tokenizer is learned end-to-end with the LLM, allowing quantization to adapt to the downstream objective. We further adopt a next-block prediction objective that models grouped discrete token blocks, mitigating context-length explosion while preserving multimodal information. Extensive experiments demonstrate state-of-the-art results on major understanding and generation benchmarks among comparable-scale unified models.
@@ 2 Design Philosophy
Our approach. We address this bottleneck with a product-quantization-based discretization scheme that substantially increases the information capacity of discrete vision tokens, yielding understanding on par with continuous-feature VLMs while remaining fully discrete and autoregressive. For the visual modality, we use a ViT-based encoder as the Tokenizer and a diffusion-based decoder as the Detokenizer, letting each component focus on its strength: the Tokenizer aligns visual signals into the shared semantic space, the LLM handles reasoning, and the Detokenizer renders high-quality images.
@@ 3 Architecture
This section begins by introducing the discrete image tokenizer, detailing how to quantize the ViT-produced patch embeddings via a product quantization mechanism. We then introduce a next-block-prediction training paradigm for the unified LLM backbone, which compresses the increased vision tokens from our multi-token tokenizer into manageable blocks, thereby preserving training and inference efficiency comparable to single-token methods. Finally, we describe the diffusion-based de-tokenizer that renders high-quality images conditioned on the LLM-generated image tokens.
@@ 3.1 Discrete Image Tokenizer
Product Quantization for Visual Tokenization. To address this bottleneck, we adopt a product quantization (PQ) (Jégou et al., 2011) strategy for visual discretization: each continuous patch feature is decomposed into $N$ parallel subspaces via learned projection matrices, and quantized independently with its own sub-codebook. The combinatorial coding capacity thus grows exponentially with $N$, substantially narrowing the information gap with continuous representations. To keep the context length unchanged for the LLM encoder, we apply sum pooling to aggregate the $N$ quantized sub-vectors back into a single composite token on the encoder side, while the decoder retains the full $N$-way discrete codes for autoregressive prediction. This asymmetric design balances representational expressivity with computational efficiency.
@@
Architecture and Training Objective. Kelix-Tok comprises two core components: (1) a NaViT-style vision encoder, initialized from Keye-VL and supporting native-resolution input, and (2) the product quantization layer described above, which decomposes each continuous patch feature into $N$ parallel discrete codes. To align the resulting discrete visual representations with the LLM text space, we attach a pre-trained lightweight language model (600M parameters) as an auxiliary decoder during training. The overall objective jointly optimizes a next-token prediction loss and a quantization loss:
@@ 3.2 Unified Model Architecture
As discussed in Section 3.1, our product quantization produces $N$ discrete codes per patch, substantially increasing the number of vision tokens compared to the original ViT output. To handle this without inflating the context length, we adopt a unified autoregressive Transformer architecture built around a Next-Block Prediction (NBP) paradigm.
@@ 5.2 Image Understanding Performance
By contrast, Kelix stands as the first discrete-tokenization unified model to achieve OCRBench performace on par with that of understanding-only VLMs. Specifically, Kelix attains an OCRBench score of 86.7, which is nearly identical to the 86.4 achieved by VLM Qwen2.5-VL (7B), while outperforming the second-best discrete-tokenization baseline X-Omni (7B) by a substantial 16.3 points, corresponding to a relative performance improvement of 23%. The result validates that Kelix’s product quantization mechanism effectively bridges the representational gap between continuous visual embeddings and discrete tokens, demonstrating that discrete visual representations are no longer a fundamental performance bottleneck for multimodal understanding even for text-rich tasks that demand fine-grained semantic preservation.
@@ 6.1 Ablation Study on Kelix-Tok
Baseline Comparison. The results in Table 6(a) reveals a critical trade-off among different visual representation schemes. We observe that, compared to the Continuous Baseline which represents a performance upper bound, the Discrete Baseline introduces a significant performance gap. This is particularly pronounced on Text-rich tasks, where its performance drops to approximately 85% of the continuous baseline (66.5 vs. 78.0 Avg.) due to the information bottleneck. Our Kelix-Tok, leveraging a PQ-based combinatorial representation, successfully achieves discretization while attaining performance comparable to its continuous counterpart.
@@
Number of Sub-Tokens. We investigate the impact of varying the number of sub-tokens $N$ by conducting experiments with $N=1,2,4,8,16$ alongside a continuous baseline, with results presented in Fig. 7. As shown in Subfigure. 7(a), the LM loss consistently decreases with increasing $N$, and the loss curves for $N\geq 8$ converge closely to the continuous baseline, indicating that sufficient discrete tokens can effectively compress continuous visual information without substantial loss. We further evaluate our training checkpoints on the OCRBench benchmark, which are highly sensitive to model performance. Subfigure. 7(b) demonstrates that OCRBench scores consistently increase with $N$. Upon training convergence, the performance of the continuous baseline is nearly matched by the top two discrete configurations ($N=8$ and $N=16$), which further validates the robustness of our PQ-based Kelix.
@@ 7.2 Unified Multimodal Models
Our Position. Our proposed Kelix follows the unified architecture route with high-level discrete semantics, but introduces a product quantization mechanism that expands the representational capacity of discrete tokens without increasing the LLM’s sequence length. This directly addresses the information loss bottleneck of discrete tokenization, aiming to close the understanding gap between discrete and continuous representations and unlock the full potential of unified multimodal modeling.